# KVComm: Enabling Efficient LLM Communication through Selective KV Sharing

### v2

@@ 1 Introduction
We evaluate KVComm on a diverse set of tasks with eight model pairs (see Section 4.1), showing that it consistently outperforms existing communication protocols while significantly reducing the data transmitted between models. In summary, our work makes three key contributions:
@@ 3.2 KV Selection Strategies
$\hat{S}_{a}^{l}=\frac{1}{HT}\sum_{h=1}^{H}\sum_{t=1}^{T}\sum_{c=1}^{|C|}a_{h,t,c}^{l},$ (1)
where $H$ is the number of attention heads, $T$ is the number of tokens in the query, $|C|$ is the number of context tokens, and $a_{h,t,c}^{l}$ is the attention weight assigned by head $h$ at layer $l$ from token $t$ to context token $c$. $\hat{S}_{a}^{l}$ is then normalized to the range $[0,1]$ across all layers to obtain the final attention importance score $S_{a}^{l}=\frac{\hat{S}_{a}^{l}-\min_{l^{\prime}}\hat{S}_{a}^{l^{\prime}}}{\max_{l^{\prime}}\hat{S}_{a}^{l^{\prime}}-\min_{l^{\prime}}\hat{S}_{a}^{l^{\prime}}}$.
@@ 3.3 Complexity Analysis
We analyze the computational complexity of our KVComm framework compared to baseline methods. Compared to the NLD (Du et al., 2023) method, our method does not require multiple decoding steps for $\mathcal{M}_{s}$, which significantly reduces the computation cost. When the number of tokens generated during debate (Du et al., 2023) is large, the computation margin of our method over NLD is on the order of $O(L(T_{s}+T_{r}+|Q|)^{2}d),$ where $T_{s}$ and $T_{r}$ are the number of tokens generated by $\mathcal{M}_{s}$ and $\mathcal{M}_{r}$ in the debate, respectively, and $|Q|$ and $d$ are the number of tokens in the query and the hidden dimension of the model, respectively. Compared to the Skyline (Section 4.1) method, our method also reduces the computation cost, especially when $M$ is small. The computation margin of our method over Skyline is on the order of O(|C|d(L(2|Q|+T)M(|Q|+T)))O(|C|d(L(2|Q|+T)-M(|Q|+T))), where $|C|$ is the number of tokens in the context, and $|T|$ is the number of tokens generated by $\mathcal{M}_{r}$.
@@ 4.1 Experimental Setup
We conduct experiments on eight different model pairs, shown in Table 5 in Section B.3. The model pairs include two instances of the same LLM and two models that are fine-tuned versions of the same base LLM. These models cover different families, including LLaMA (Dubey et al., 2024), Qwen (Qwen et al., 2024), and Falcon (Almazrouei et al., 2023).
@@ 4.5 Attention Distribution Analysis
We validate hypothesis H2 in Section 3.2 by selecting layers with different attention importance scores for communication. We select $9$ layers with different levels of attention importance scores, and test the communication performance with Llama-3.2-3B model. The results are shown in Figure 8. We can find that selecting layers with higher scores can achieve better performance, while selecting layers with lower scores can diminish the performance. This validates hypothesis H2 that layers with higher attention importance scores are more effective for communication.
@@ 4.6 System Efficiency
Mathematically, we have shown in Section 3.3 that KVComm can reduce the computation cost compared to Skyline. We validate this through experiments on the Llama-3.2-3B model pair with Tipsheets and MultiFieldQA-en datasets. We report the relative FLOPs of KVComm and Skyline over AC in Figure 8. NLD and CIPHER are not included since they require multiple decoding steps for $\mathcal{M}_{s}$, which makes the computation cost significantly higher than AC. We can find that KVComm has a significant computation advantage over Skyline, especially when selecting fewer layers for communication. This demonstrates the efficiency of our KVComm framework in enabling effective communication with reduced computational overhead by 2.5x to 6x.
In addition to FLOPs, we also report the memory consumption among methods. KVComm similarly shows a substantial memory advantage over Skyline, as the reduced number of communicated layers not only lowers computation but also alleviates memory pressure. On Tipsheets, KVComm uses 23% to 73% less memory than Skyline. KVComm This further highlights the efficiency of our KVComm framework in achieving lightweight inter-model communication.
@@ B.3 Model Pairs
We conduct experiments on eight different model pairs, shown in Table 5. The first four pairs consist of the same LLMs, while the last four pairs consist of models that are fine-tuned on the same base LLM.

### v3

@@ 1 Introduction
We evaluate KVComm on a diverse set of tasks with nine model pairs (see Section 4.1), showing that it consistently outperforms existing communication protocols while significantly reducing the data transmitted between models. In summary, our work makes three key contributions:
@@ 3.2 KV Selection Strategies
$\hat{S}_{a}^{l}=\frac{1}{H|Q|}\sum_{h=1}^{H}\sum_{q=1}^{|Q|}\sum_{c=1}^{|C|}a_{h,q,c}^{l},$ (1)
where $H$ is the number of attention heads, $|Q|$ is the number of tokens in the query, $|C|$ is the number of context tokens, and $a_{h,q,c}^{l}$ is the attention weight assigned by head $h$ at layer $l$ from token $q$ to context token $c$. $\hat{S}_{a}^{l}$ is then normalized to the range $[0,1]$ across all layers to obtain the final attention importance score $S_{a}^{l}=\frac{\hat{S}_{a}^{l}-\min_{l^{\prime}}\hat{S}_{a}^{l^{\prime}}}{\max_{l^{\prime}}\hat{S}_{a}^{l^{\prime}}-\min_{l^{\prime}}\hat{S}_{a}^{l^{\prime}}}$.
@@ 3.3 Complexity Analysis
We analyze the computational complexity of our KVComm framework compared to baseline methods. Compared to the NLD (Du et al., 2023) method, our method does not require multiple decoding steps for $\mathcal{M}_{s}$, which significantly reduces the computation cost. When the number of tokens generated during debate (Du et al., 2023) is large, the computation margin of our method over NLD is on the order of $O(L(T_{s}+T_{r}+|Q|)^{2}d),$ where $T_{s}$ and $T_{r}$ are the number of tokens generated by $\mathcal{M}_{s}$ and $\mathcal{M}_{r}$ in the debate, respectively, and $|Q|$ and $d$ are the number of tokens in the query and the hidden dimension of the model, respectively. Compared to the Skyline (Section 4.1) method, our method also reduces the computation cost, especially when $M$ is small. The computation margin of our method over Skyline is on the order of $O(|C|d(L(2|Q|+T_{r})-M(|Q|+T_{r})))$, where $|C|$ is the number of tokens in the context.
@@ 4.1 Experimental Setup
We conduct experiments on nine different model pairs, shown in Table 5 in Section B.3. The model pairs include two instances of the same LLM and two models that are fine-tuned versions of the same base LLM. These models cover different families, including LLaMA (Dubey et al., 2024), Qwen (Qwen et al., 2024), and Falcon (Almazrouei et al., 2023).
@@ 4.5 Attention Distribution Analysis
We validate hypothesis H2 in Section 3.2 by selecting layers with different attention importance scores for communication. We select $9$ layers with different levels of attention importance scores, and test the communication performance with Llama-3.2-3B model. The results are shown in Figure 8. We find that selecting layers with higher scores can achieve better performance, while selecting layers with lower scores can diminish the performance. This validates hypothesis H2 that layers with higher attention importance scores are more effective for communication.
@@ 4.6 System Efficiency
Mathematically, we have shown in Section 3.3 that KVComm can reduce the computation cost compared to Skyline. We validate this through experiments on the Llama-3.2-3B model pair with Tipsheets and MultiFieldQA-en datasets. We report the relative FLOPs of KVComm and Skyline over AC in Figure 8. NLD and CIPHER are not included since they require multiple decoding steps for $\mathcal{M}_{s}$, which makes the computation cost significantly higher than AC. We find that KVComm has a significant computation advantage over Skyline, especially when selecting fewer layers for communication. This demonstrates the efficiency of our KVComm framework in enabling effective communication with reduced computational overhead by 2.5x to 6x.
In addition to FLOPs, we also report the memory consumption among methods. KVComm similarly shows a substantial memory advantage over Skyline, as the reduced number of communicated layers not only lowers computation but also alleviates memory pressure. On Tipsheets, KVComm uses 23% to 73% less memory than Skyline. This further highlights the efficiency of our framework in achieving lightweight inter-model communication.
@@ B.3 Model Pairs
We conduct experiments on nine different model pairs, shown in Table 5. The first four pairs consist of the same LLMs, while the last five pairs consist of models that are fine-tuned on the same base LLM.